The Town of Bedford Falls approved a General Fund operating budget for the fiscal year beginning on July 1. The budget provides for estimated revenues of $2,539,000 as follows: property taxes, $1,844,000; licenses and permits, $315,000; fines and forfeits, $215,000; and intergovernmental (state grants), $165,000. The budget approved appropriations of $2,496,000 as follows: General Government, $465,000; Public Safety, $1,530,000; Public Works, $315,000; Culture and Recreation, $143,000; and Miscellaneous, $43,000.

Required

  1. a. & b. Prepare the journal entry (or entries), to record the Town of Bedford Falls’s General Fund operating budget on July 1, the beginning of the Town’s fiscal year and also record the following transactions that occurred during the month of July. Also show entries in the subsidiary ledger accounts, to record the Town of Bedford Falls’s General Fund operating budget on July 1, the beginning of the Town’s fiscal year. (If no entry is required for a transaction/event, select "No Journal Entry Required" in the first account field.)

  1. Revenues were collected in cash amounting to $48,000 for licenses and permits and $20,500 for fines and forfeits.
  2. Supplies were ordered by the following functions in early July at the estimated costs shown:
General Government $ 9,100
Public Safety 14,700
Public Works 7,800
Culture and Recreation 5,900
Miscellaneous 2,600
Total $ 40,100
  1. During July, supplies were received at the actual costs shown below and were paid in cash. General Government, Culture and Recreation, and Miscellaneous received all supplies ordered. Public Safety and Public Works received part of the supplies ordered earlier in the month at estimated costs of $12,400 and $7,600, respectively.
Actual Cost Estimated Cost
General Government $ 9,000 $ 9,100
Public Safety 12,500 12,400
Public Works 7,800 7,600
Culture and Recreation 5,800 5,900
Miscellaneous 2,600 2,600
Total $ 37,700 $ 37,600
  1. Calculate the amount of budgeted but unrealized revenues in total and from each source as of July 31. (Input all amounts as positive values.)
  1. Calculate the amount of available appropriation in total and for each function as of July 31.(Input all amounts as positive values.)
Answer #1

Please find below spreadsheet useful to get desired results: -

A

General Ledger Subsidiary Ledger 1 a) Debits Credits Debits Credits 2 3 ESTIMATED REVENUES 2,539,000 BUDGETARY FUND 4 43,000

B.

Н R Subsidiary Ledger General Ledger Debits Credits Debits Subsidiary Ledger b) General Ledger Debits Credits Debits Credits

C and D

CALCULATION OF BUDGETED BUT UNREALIZED REVENUES AS OF JULY 31, 2016 UNREALIZED 3 SOURCE BUDGETED ACTUAL REVENUE 5 PROPERTY TA

Note: -

Encumbrances

Public safety = 14,700-12,400 = 2,300

Public works = 7,800-7,600 = 200